A 0.5-L unbuffered solution needs the pH adjusted from 6 to 3. How many microliters of a 2 molar HCl solution need to be added to adjust the pH?

Report your answer to four decimal places.

For this question:

  • Assume the volume of HCl being added is negligible and the total volume of the solution does not change.
  • The volume of the solution is *not* 1 L like it was in the practice pools.
  • The question requests your answer be reported in μL (microliters)!!
Answer #1

We know that: pH = -log([H+])

From this: [H+] = 10-pH

Initial [H+] = 10-6 M

Final [H+] = 10-3 M

Change in [H+] = 10-3 - 10-6 = 9.99 * 10-5 M

Moles of H+ = [H+] * Volume of solution = 9.99 * 10-5 * 0.5 = 5 * 10-5 moles

Molarity of HCl * Volume of HCl = moles of H+

2 * V = 5 * 10-5

Solving, V = 2.5 * 10-5 = 25 microliters